OpenStudy (anonymous):

Frank kicks a soccer ball off the ground and in the air with an initial velocity of 30 feet per second. Using the formula H(t) = −16t2 + vt + s, what is the maximum height the soccer ball reaches?

OpenStudy (anonymous):

a. 14.1 feet b. 13.7 feet c. 13.2 feet d. 15.2 feet

OpenStudy (phi):

they tell you "Frank kicks a soccer ball off the ground " which means s in the equation is 0 the way you know this, is to say, at time t=0 (at the start) H(0)= -16*0*0+ v*0 + s or H(0)= s if the ball is on the ground at the start, its height is 0, os H(0)=0 = s they tell you "initial velocity of 30 feet per second. " that means v=30 your equation is H(t) = -16 t^2 + 30 t that is a parabola. You need to find the vertex (highest point) of this vertex

OpenStudy (anonymous):

Oh okay. Thanks!

OpenStudy (phi):

see http://hotmath.com/hotmath_help/topics/vertex-of-a-parabola.html for how to find the t value of the vertex. use that t value to find the height

OpenStudy (anonymous):

Got it!!! :) thanks<3